OpenStudy (anonymous):

find 2 positive real #s whose sum is 13, if the sum of their squares is a min

OpenStudy (anonymous):

sum of their squares is ....................

OpenStudy (anonymous):

a minimum

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6.5 again?

OpenStudy (anonymous):

lools yeah nn can u do it using a quadratic equation .. step by step

OpenStudy (anonymous):

a + b = 13 a^2 + b^2 is minimum. => a^2 + b^2 = 169 - 2ab Also ab<169/4 Therefore for Minimum a^2 + b^2 we take maximum ab. Which => a^2 + b^2 = 169/2 And (a - 13)^2 = 169/2 - a^2

OpenStudy (anonymous):

Solve the quadratic to get two real values of x. Oh and ab, 169/4 by AM>GM a+b/2 > (ab)^1/2

OpenStudy (anonymous):

* ab <(equalto) 169/4

OpenStudy (anonymous):

i guess i kind get it.. thnxx

OpenStudy (anonymous):

Haha. Anytime :)

OpenStudy (anonymous):

=D